OpenStudy (anonymous):

PLEASE HELP!! I WILL GIVE A MEDAL!!!!! You have $59.95 in your wallet and want to buy some new CDs. If the CDs are $11.99 each, what number of CDs, x, can you buy? Write and solve an inequality.

OpenStudy (anonymous):

Is this correct? \[11.99x \le 59.95\] \[\frac{ 11.99x }{ 11.99 } \le \frac{ 59.95 }{ 11.99 }\] \[x \le5\]

OpenStudy (anonymous):

@Compassionate @ganeshie8 @goformit100

OpenStudy (taylor<3srin):

This is basically just dividing 59.99 by 11.99.

OpenStudy (anonymous):

I know but I have to write it in an inequality and solve it. Do you think the inequality I made above is correct?

OpenStudy (taylor<3srin):

It should be okay